# Predicting and Understanding Turn-Taking Behavior in Open-Ended Group Activities in Virtual Reality

### v1

@@ 4.3. Feature Interpretation Analysis
We conducted additional analyses on how features with high importance are related to model predictions. Specifically, we interpreted features by their partial dependence, which is a commonly used technique used to interpret machine learning models such as random forests (Furlanello et al., 2003; Auret and Aldrich, 2012) and gradient boosting classifiers (Natekin and Knoll, 2013; Blagus and Lusa, 2017). By definition, partial dependence measures the marginal effect on model prediction through varying the values to a feature and calculating the average probability estimates after this procedure. We can then interpret the resulting plot of partial dependence and feature value as the relationship between that feature and model predictions. In our setup, we used models trained on the entire dataset and varied each feature between its [0.05, 0.95] percentile.
We selected features within each feature group that had relatively high feature importance. We did not interpret prior speech sequence features as past research has suggested the need to consider the interactions between prior speakers when modeling speech behaviors (Parker, 1988). In aggregate, we analyzed 56 features. For speech-related features, we examined 2 features: the user turn count and user previous turn time. There were 6 features related to individual- and group differences, namely the Big-5 personalities and group size. 12 features were related to the group relationship, specifically those capturing the group’s direct gaze towards the main user and reference user, respectively. The remaining 36 were egocentric motion features, namely those describing the head position in the y-axis, rotation in the yaw axis, and left hand position in the y-axis for the main and reference users, respectively.
![Refer to caption](extracted/5634250/figures/Partial-Dependencies-v3.png)
![Refer to caption](extracted/5634250/figures/Partial-Dependencies-contd-v3.png)
@@ 4.3.1. Speech-Related Features
Seen in Figure 4a, varying the amount of time and number of turns to the main user’s previous turn revealed large effects on the probability estimates of data samples. Specifically, the results suggested that listeners (i.e., main users) who had spoken more recently are more likely to engage in turn-taking behaviors and be the next speaker (i.e., predicting turn taking vs. continuing speech and the next speaker). The models also predicted that listeners are more likely to speak at a time closer to their previous speech activity. Our results on the user turn count feature revealed for the next speaker prediction task, listeners whose last speaking event occurred at earlier turns are less likely to become the next speaker.
4.3.2. Individual- and Group- Level Difference
As shown in Figure 4b, we first note that for predicting the next speaker and differentiating between turn taking and continuing speech, listeners with higher values of extraversion are more likely to be the next speaker. There was little variation in probabilities estimates when varying main user personalities for predicting the timing of turn-taking behaviors. This suggests that listeners with different levels of Big-5 personalities do not exhibit different patterns in when they decide to speak. There were weaker but noticeable relationships between the listener’s consciousness, neuroticism, and openness and probabilities estimates. For example, for the next speaker prediction task, the model predicted that listeners reporting a lower level of openness were less likely to be the next speaker.
@@ 4.3.3. Group Relationship
The most visually salient relationship in the partial dependence plots for examined features regarding group relationships (Figure 4c) were those related to the group’s direct gaze towards the previous speaker. Notably, the models predicted that a new speaker is more likely to take over a previous turn in moments when listeners are looking away from the previous speaker (i.e., larger minimum and average value for the direct gaze from the group to the reference user). We note a similar but weaker trend when predicting the timing of a turn taking.
The partial dependence plots for the group’s direct gaze towards a listener revealed the opposite directionality between prediction estimates and feature values, where a listener is predicted to be more likely to take over a previous turn when the remaining group is looking more directly at them (i.e., small maximum and average value for the direct gaze from the group to the main user). This pattern is, however, less pronounced than the greater variability in probability estimates exhibited for direct gaze measurements from the group to the previous speaker.
@@ 4.3.4. Egocentric Motion
Figure 5 presents our findings on selected egocentric motion features. As shown in Figure 5a, the feature with the greatest variation in partial dependence for the left hand y-axis position was that of the average velocity for both the listener and previous speaker. Most notably, the model used for differentiating between turn taking and continuing speech predicted that previous speakers who move their left hand upward in the vertical axis (i.e., large positive values in the average y-axis velocity) are more likely to continue speaking. In the case for listeners, models for all three tasks predicted that listeners are more likely to speak in moments right after they move their left hand upward (i.e., large positive values in the average y-axis velocity).
For egocentric features related to head y-axis position, the salient trends were related to velocities (Figure 5b). To start, the model for the timing of the turn taking task predicted that a listener is more likely to begin speaking if the previous speaker’s head had on average not moved a substantial amount in either direction in the vertical axis. This was, however, not the case for the models for the other two tasks. On the contrary, all three models predicted that a listener is more likely to begin speaking if their head had on average moved upward in the y-axis (i.e., large positive values for the average y-axis velocity). Relatedly, a listener is predicted to also be more likely to speak if their head had at any point in the sampled window duration moved quickly upward in the vertical axis (i.e., large positive values for the maximum y-axis velocity).
Finally, as seen in Figure 5c, for features related to head pitch rotation, we found distinctive patterns in maximum and minimum velocities of the pitch rotations for the main and reference users. We interpret this as the models finding the extremes of an individual’s head pitch rotation velocities across the 1-second window more informative than their averages. There were no noticeable trends in partial dependence for the previous speaker’s head pitch rotation on the next speaker prediction task. Models for the remaining two tasks predicted that a listener is less likely to speak if the previous speaker has at any point extensively rotated their head upward, which in our coordinate system maps to small negative values for the minimum pitch velocity. The models also predicted that a listener is less likely to speak if the previous speaker has at any moment within the 1-second window extensively rotated their head downward (i.e., large positive values for the maximum pitch velocity). The visual trends in partial dependence on the listener’s head pitch rotation were opposite to that of the previous speaker. Namely, all three models predicted a listener is more likely to speak when they have engaged in substantial downward head rotation (i.e., large positive values for the maximum pitch velocity) or upward head rotation (i.e., small negative values for the minimum pitch velocity).
@@ 5. Discussion
5.1. Practical Implications of Real-time Turn-Taking Behavior Prediction
@@
With regards to RQ1, we found that the gradient boosting classifiers achieved consistently the best performance on all three tasks across different performance metrics. This suggests that the relationship between the extracted features and the outcome variable is non-linear and that interactions between features is likely relevant to the predicted behavior. Specifically, the models obtained accuracies of 0.75–0.78 AUC on the tasks of identifying the new speaker and distinguishing between moments prior to turn transitions and continuing speech, and had poorer performance on distinguishing when a new user will begin speaking with accuracies of 0.71–0.72 AUC. These accuracies are considerably higher than prediction by chance and comparable to prior research (Chen et al., 2024). In extending our methods into tools and applications that leverage predictions of turn-taking behavior, practitioners can consider prioritizing features that can benefit from real-time knowledge of who the next speaker is, as opposed to when exactly they will begin speaking.
@@
To answer RQ2, we showed that predictive performance is robust over time and across unseen groups, with accuracies for evaluations based on unseen groups being slightly lower but comparable to the benchmark metrics. Considering also the diverse group activities and the mobility allowed in virtual spaces in our examined group sessions, these results suggest that practitioners can expect robustness across time and unseen activities and groups when training predictive models using similar approaches and deploying them into real-world social VR platforms. One possible explanation for the robust models is the high number of groups and individuals and wide range of activities already seen during training, which likely allowed the models to extrapolate generalizable insights about the prediction tasks.
@@ 5.2. Theoretical Implications of Turn-Taking Behaviors in VR
Our analysis on feature importance and interpretation (RQ3) revealed that features on listener’s personalities, group size, speech-related behavior, group gaze towards the listener and previous speaker, and the users’ left hand y-axis position and head y-axis position and pitch rotation were prominently influencing prediction accuracies and probability estimates.
@@
The results on individual- and group- level differences extend past works demonstrating that personalities such as extraversion, agreeableness, and neuroticism are correlated with user’s verbal behaviors (Lepri et al., 2010; Aran and Gatica-Perez, 2013). For example, we found that listeners who are more extroverted are predicted to be more likely to be the next speaker, which corroborated past findings that extraversion is related to speaking time (Lepri et al., 2010) and speaking turns (Aran and Gatica-Perez, 2013). Our analysis on group size suggested intuitively that a listener in a three-person group is more likely to be the next speaker compared to a listener in a four-person group. Though we focused on interpreting features individually, it is also possible that the models used group size in combination with other motion-related features for their predictions, as group sizes can be related to different head orientation patterns (Hadley et al., 2021). These findings highlight the need to consider individual and group characteristics when predicting real-time social dynamics in VR and instantiating virtual agents with distinct personas and natural behaviors. One possible approach is to continue leveraging the affordances of VR to study large and diverse demographics with high variances in individual differences, and doing so in controlled contextual and social settings (Wang and Bailenson, [n. d.]).
@@
Speech-related features exhibited high importance and salient effects on probability estimates. Specifically, our models predicted that those who more recently spoke, either in terms of speaking turns and time, are more likely to be the next speaker. While we did not analyze speech sequence features given the likely interactions between turn indices (Parker, 1988), it also provided important signals to predicting speech behaviors, in particular who the next speaker will be.
@@
Our results revealed that group gaze behavior is useful in distinguishing between moments of turn transition and continuing speech, with the individual that the group is looking more directly towards predicted to be more likely the upcoming speaker. Our findings extend past insights that gaze behavior is useful in modeling speech behaviors (Jokinen et al., 2013), classroom discourse (Stark et al., 2024), and visual attention (Jokinen et al., 2009) by demonstrating that the social group’s direct gaze towards individual users is still predictive of the next speaker even when VR groups are not constrained or assigned to fixed positions. While features related to direct gaze between the listener and previous speaker did not exhibit high importance, it is possible that these features may offer key insights on speech behaviors when sampled at moments other than those immediately preceding a turn transition, for example during speech events or at moments following the end of a previous speaking turn. Finally, our models predicted that listeners who had engaged in substantial rotation in head pitch (e.g., nodding, abruptly looking up or down) and had moved their heads and left hands vertically upward exhibited more speaking intentions. An additional analysis detailed in Appendix C on head pitch revealed that the combination of high instantaneous speeds in both the upward and downward directions can be further indicative of speaking intentions.
@@
We also found that the previous speaker’s left hand y-axis position and the head y-axis position and pitch rotation also influenced performance, though their effects on the probability estimates are less consistent across tasks. We believe that this could be due to the differences in role the previous speaker plays in the three tasks. For example, while the previous speaker’s egocentric motion may be relevant when predicting whether they will continue to speak, it is not informative for predicting who out of the remaining listeners will begin speaking next as the extracted features for the previous speaker are the same for all listeners. One explanation for the finding on left hand motion is that they are associated with the users raising the left hand to view the menu, tablet, or muting and unmuting button located on the virtual hand. It is also worth noting that we empirically observed few instances of users muting and unmuting themselves, likely due to the small group sizes and the open-ended group activities prompting more natural conversations. These findings corroborate past research demonstrating the predictive capabilities of motion tracking data for speech behavior (Ishii et al., 2017; Chen et al., 2024), and contribute additional insights on how egocentric motion features are related to the predicted behaviors.

### v2

@@ 4.3. Feature Interpretation Analysis
To study RQ3, we analyzed how features with high importance are related to predictions of the best performing models (i.e., gradient boosting classifiers). We selected features within each feature group with high feature importance based on MDA. Notably, we included features deemed insignificant in the Wald-tests in Section 4.2.2 for two reasons. First, we interpreted the gradient boosting classifiers and not the logistic regressions. Using MDA from the gradient boosting classifers allowed us to capture insights from features that may lack significance under the assumptions of a linear model but still contribute non-linearly. Additionally, logistic regressions are sensitive to collinearity and can complicate assessments of feature significance (Dormann et al., 2013). Another issue is an overemphasis on features with small effects given a large sample (Sullivan and Feinn, 2012). By leveraging MDA instead, we selected features that best reflected their contribution to the best performing models.
In aggregate, we analyzed 56 constructs. For speech-related features, we examined 2 constructs: the user turn count and user previous turn time. There were 6 constructs related to individual and group differences: main user personalities and group size. 12 were related to the group relationship, specifically those capturing the group’s direct visual attention towards the main user and reference user. The remaining 36 were egocentric motion constructs, namely those describing the head position in the y-axis, rotation in the yaw axis, and left hand position in the y-axis for the main and reference users. We did not interpret individual prior speech sequence features as research has suggested the need to consider the interactions between prior speakers (Parker, 1988).
We interpreted features using partial dependence, a common technique used to interpret machine learning models such as random forests (Furlanello et al., 2003; Auret and Aldrich, 2012) and gradient boosting classifiers (Natekin and Knoll, 2013; Blagus and Lusa, 2017). By definition, partial dependence measures the marginal effect on model prediction through varying the values to a feature and calculating the average probability estimates after this procedure. In our setup, we chose the gradient boosting classifiers trained on the entire dataset and varied each feature between its [0.05, 0.95] percentile.
@@ 4.3.1. Speech-Related Features
Seen in Figure 4a, varying the amount of time and number of turns to the main user’s previous turn revealed large effects on the probability estimates. Specifically, the results suggested that listeners who had spoken more recently are more likely to engage in turn-taking behaviors and be the next speaker (i.e., predicting turn taking vs. continuing speech and the next speaker). The models also predicted that listeners are more likely to speak at a time closer to their previous speech activity. Our results on the user turn count feature revealed for the next speaker prediction task that listeners whose last speaking event occurred at earlier turns are less likely to become the next speaker.
4.3.2. Individual- and Group- Level Differences
Shown in Figure 4b, for predicting the next speaker and differentiating between turn taking and continuing speech, listeners with higher values of extraversion are more likely to be the next speaker. There was little variation in probabilities estimates when varying main user personalities for predicting the timing of turn-taking. This suggests that listeners with different levels of Big-5 personalities do not exhibit different patterns in when they decide to speak. There were weaker but noticeable relationships between the probabilities estimates and the listener’s consciousness, neuroticism, and openness. For example, for the next speaker prediction task, the model predicted that listeners reporting a lower level of openness were less likely to be the next speaker.
@@ 4.3.3. Group Relationship
The most salient relationship in the partial dependence plots regarding group relationships (Figure 4c) were those related to the group’s direct visual attention towards the previous speaker. Notably, the models predicted that a new speaker is more likely to take over a previous turn when listeners are looking away from the previous speaker (i.e., larger minimum and average value for the direct visual attention angle from the group to the reference user). There is a similar but weaker trend when predicting the timing of turn taking.
The partial dependence plots for the group’s direct visual attention towards a listener revealed the opposite directionality between prediction estimates and feature values, where a listener is predicted to be more likely to take over a turn when the remaining group is looking more directly at them (i.e., small maximum and average value for the direct visual attention angle from the group to the main user). This pattern is, however, less pronounced than the greater variability in probability estimates exhibited for direct visual attention measurements from the group to the previous speaker.
@@ 4.3.4. Egocentric Motion
Figure 5 presents our findings on egocentric motion features. Shown in Figure 5a, the feature with the greatest variation in partial dependence for the left hand y-axis position was the average velocities for both the listener and previous speaker. Notably, the model used for differentiating between turn taking and continuing speech predicted that previous speakers who move their left hand upward in the vertical axis (i.e., large positive values in the average y-axis velocity) are more likely to continue speaking. For listeners, models for all three tasks predicted that listeners are more likely to speak in moments right after they move their left hand upward (i.e., large positive values in the average y-axis velocity).
For egocentric features related to head y-axis position, the salient trends were related to velocities (Figure 5b). To start, the model for the turn-taking timing task predicted that a listener is more likely to begin speaking if the previous speaker’s head had on average not moved a substantial amount in either direction in the vertical axis. This was not the case for models for the other two tasks. On the contrary, all three models predicted that a listener is more likely to begin speaking if their head had on average moved upward in the y-axis (i.e., large positive values for the average y-axis velocity). Relatedly, a listener is predicted to more likely speak if their head had at any point in the sampled window moved quickly upward in the vertical axis (i.e., large positive values for the maximum y-axis velocity).
Finally, for features related to head pitch rotation, we found distinctive patterns in maximum and minimum velocities of the pitch rotations for the main and reference users (Figure 5c). We interpret this as the models finding the extremes of an individual’s head pitch rotation velocities across the 1-second window more informative than their averages. There were no noticeable trends in partial dependence for the previous speaker’s head pitch rotation on the next speaker prediction task. Models for the remaining two tasks predicted that a listener is less likely to speak if the previous speaker has at any point extensively rotated their head upward, which corresponds to small negative values for the minimum pitch velocity. The models also predicted that a listener is less likely to speak if the previous speaker has at any moment within the 1-second window extensively rotated their head downward (i.e., large positive values for the maximum pitch velocity). The trends in partial dependence on the listener’s head pitch rotation were opposite to that of the previous speaker. Namely, all three models predicted a listener is more likely to speak when they have engaged in substantial downward head rotation (i.e., large positive values for the maximum pitch velocity) or upward head rotation (i.e., small negative values for the minimum pitch velocity).
@@ 5. Discussion
5.1. Practical Implications of Turn-Taking Prediction
5.1.1. Model Selection and Performance
@@
With regards to RQ1, we found that the gradient boosting classifiers achieved consistently the best performance on all tasks across performance metrics. This suggests that the relationship between the features and outcome variable is non-linear and that interactions between features are likely relevant for prediction. Specifically, the models obtained accuracies of 0.75–0.78 AUC on the tasks of identifying the new speaker and distinguishing between turn transitions and continuing speech, and had poorer performance on predicting when a new user will begin speaking with accuracies of 0.71–0.72 AUC. These accuracies are considerably higher than prediction by chance, are considered to have excellent discrimination between positive and negative samples (Yang and Berdine, 2017), and comparable to prior research (Chen et al., 2024).
As AUC represents the probability that a random positive sample is rated higher than a random negative sample (Bradley, 1997), there remains room for prediction improvements, albeit limited by the complexity inherent to social dynamics and VR (Degutyte and Astell, 2021; Maloney et al., 2020). To further explore the practical implications of turn-taking predictions, we present an additional analysis in Appendix E. In it, we consider the task of predicting continuing speech and turn-taking events and extended the binary classification problem into a multiclass one. This analysis yielded comparable findings and found that three-person groups yielded higher prediction accuracy than four-person ones. One possible explanation for the lower performance in larger groups is the greater unpredictability in larger-group social dynamics. While there were more four-person discussions than three-person ones in our dataset, greater complexity in social dynamics may warrant more data for larger group sizes. When modeling and deploying comparable models in real-world settings, our findings here suggest that practitioners should sample diverse group sizes during data collection, exercise caution when predicting turn-taking behaviors for groups whose sizes differ from the training data, and not assume higher prediction accuracies in smaller groups will translate directly to larger groups.
Additionally, our model accuracies suggest that near-term uses cases should prioritize applications that either 1) are minimally impacted by potential false labels or 2) build off insights from feature importance and interpretation. Given the lower performance in predicting turn taking timing, practitioners should build tools that leverage predictions for who the next speaker is, as opposed to when they will begin speaking.
@@
Importantly, overall accuracy alone does not capture the full picture to model performance, for generalizability is critical for practical deployments. In our work, we showed that predictive performance is robust over time and across unseen groups and activities (RQ2), with accuracies for evaluations based on unseen groups being slightly lower but comparable to benchmark metrics. Considering also the diverse group activities and virtual mobility allowed in our dataset, these results suggest that practitioners can expect robustness across time and unseen activities and groups when training and deploying predictive models using similar approaches. One possible explanation for the robustness is the high number of groups and wide range of activities seen during training allowing models to extrapolate generalizable insights.
5.1.2. Application of Findings
Our findings offer insights for practical applications. For one, for educators and organizers facilitating group discussions, having a tool for predicting next speakers can help with moderation (e.g., smoother transition between speakers). As these facilitators may need to multi-task and monitor multiple sessions (Sato et al., 2023), such tools can reduce the cognitive load for interpreting the nonverbal behavior of multiple social scenes concurrently. Another use case is in facilitating interactions with users joining at different levels of immersion (e.g., audio only, full body tracking). For example, model that predict and notify audio-only users of the speaking intentions of immersive users can help them better navigate social scenarios. Conversely, systems can model nonverbal behaviors from audio-only users and notify fully-immersed users of possible speaking intentions of audio-only users, thereby addressing the asymmetry in immersion.
Another application is in training user awareness in detecting turn-taking behaviors. One option is to visually guide users to notice predictive nonverbal features before speaking turns, similar to immersive training systems (Huang et al., 2021; Brough et al., 2007). Such training can be beneficial to populations who struggle socially (Freeman et al., 2002; Hall et al., 1999) and professionals in domains such as education and healthcare. Our features analysis on individual and group differences also offers implications on group composition. Namely, aligning with works on social dynamics and individual differences (Ramsay, 1968; Beukeboom et al., 2013; Aran and Gatica-Perez, 2013; Lepri et al., 2010), practitioners interested in shaping turn-taking behaviors can vary group size and leverage individuals characteristics to form the “ideal social group.”
Finally, our work sheds light on virtual agents instantiation. Conversational agents struggle with interrupting at appropriate times (Maier et al., 2022), and we foresee future systems leveraging turn-taking predictions to decide whether and when to interject in natural and non-intrusive ways. Possible application of virtual agents include teachers in classrooms, moderators in focus groups and conferences, and facilitators in support groups.
@@ 5.2. Theoretical Implications of Turn-Taking Behaviors in VR
Understanding how features predict turn-taking behaviors offer practical benefits (e.g., smaller models with curated features) and theoretical implications (RQ3). Here, we expand on the latter, drawing parallels to past literature and supplementing them with post-hoc analysis.
@@
The results on individual and group differences extend works demonstrating that personalities such as extraversion, agreeableness, and neuroticism are correlated with user’s verbal behaviors (Lepri et al., 2010; Aran and Gatica-Perez, 2013). As noted in Appendix C, features related to personalities and group sizes significantly contributed to logistic regression models in at least one of the three tasks. Additionally, we found that more extroverted listeners are predicted to more likely be the next speaker, corroborating past findings that extraversion is related to speaking time (Lepri et al., 2010) and speaking turns (Aran and Gatica-Perez, 2013). Our analysis on group size suggested that listeners in three-person groups are more likely to be the next speaker compared to listeners in four-person groups. Though we focused on individual feature interpretation, it is also possible that models used group size in combination with motion-related features for their predictions, as group size can be related to head orientations (Hadley et al., 2021). Another related theoretical thread to consider regarding group size is how differences in group composition (e.g., personality, demographics) may affect social dynamics and psychological processes differently. For example, larger groups, compared to smaller ones, tend to report lower levels of group identity and greater role differentiation in performing group tasks, and consequently require more control from leadership to coordinate efforts across group members (Hare, 1981). As such, with many of the collaborative activities participants engaged in within our dataset (see Appendix A for the list of discussion and group activities), we imagine more drastic and fundamental differences in group behaviors, both verbally and nonverbally, across group that greatly differ in size. While we studied three to four person social groups, just how comparable and similar are turn-taking behaviors between groups examined in this paper and those from much larger groups (e.g., over 10 people)? These findings and considerations highlight the need to consider individual and group characteristics when predicting social dynamics and instantiating virtual agents with distinct personas and natural behaviors. One possible approach is to continue leveraging VR to study large and diverse demographics with high variances in individual differences, and doing so in controlled settings (Wang and Bailenson, [n. d.]).
@@
Speech-related features exhibited high importance and salient effects on probability estimates for the gradient boosting classifiers, and were also significant predictors for the logistic regressions. Furthermore, our models predicted that those who more recently spoke, both in terms of speaking turns and time, are more likely to be the next speaker. While we did not analyze speech sequence features given the likely interactions between turn indices (Parker, 1988), it provided important signals to predicting speech behaviors, in particular to next speaker prediction.
@@
Our results revealed that group visual attention behavior is useful in distinguishing between turn transition and continuing speech, with the individual that the group is looking more directly towards predicted to be more likely the upcoming speaker. These findings extend past insights that visual attention measurements are useful in modeling speech behaviors (Jokinen et al., 2013, 2009) and classroom discourse (Stark et al., 2024) by demonstrating that the social group’s direct visual attention towards individuals is still predictive of the next speaker even when VR groups are not constrained to fixed virtual positions. While features related to direct visual attention between the listener and previous speaker were significant in predicting the next speaker in logistic regressions, they did not exhibit high importance for gradient boosting models. Nevertheless, these features could still offer insights on speech behaviors when sampled at other moments, for example during speech events or immediately after speaking turns.
@@
Our models predicted that listeners who engaged in substantial rotation in head pitch (e.g., nodding, abruptly looking up or down) and moved their heads and left hands vertically upward exhibited more speaking intentions. Our feature significance results largely echoed the importance of these features. An additional analysis detailed in Appendix F on head pitch revealed that the combination of high instantaneous speeds in both the upward and downward directions can be further indicative of speaking intentions. We also found that the previous speaker’s left hand y-axis position and the head y-axis position and pitch rotation influenced performance, though their feature significance and effects on probability estimates are less consistent across tasks. Of note, features related to the previous speaker’s left hand y-axis position were not significant in the logistic regressions. Possible explanations include their potential interactions with other features and non-linear contributions to the predictions. The inconsistency across tasks could be due to the different roles the previous speaker plays in the tasks. For example, while the previous speaker’s egocentric motion could help predict whether they will continue to speak, it is not informative for predicting which listener will speak next. One explanation for the finding on left-hand motion is its association with users raising their hands to access the menu, tablet, and audio button. These findings corroborate past research demonstrating the predictive capabilities of tracking data for modeling speech behavior (Ishii et al., 2017; Chen et al., 2024), and contribute insights on how egocentric motion is related to the predicted behaviors.
Importantly, though our findings align with prior literature, feature interpretation can differ across virtual context. Similar to how social factors such as group size and personality influenced our model predictions, virtual context such as room size can also impact nonverbal behaviors (Han et al., 2023, 2024). Platform-specific characteristics such as the location of virtual menus could influence body motion and visual attention preceding speech events. Though our dataset varied group size and spatial context, model performance and interpretation may still differ when evaluated on novel social settings.